Metronidazole is used in the treatment of anaerobic protozoal infections. Although often administered orally, it can also be delivered by means of an IV infusion. The loading dose is 15 mg per kilogram of body mass and is infused over one hour. The maintenance dose is 7.5 mg per kilogram of body mass, is also infused over an hour, and is administered six hours after the loading dose. If the concentration of the metronidazole solution is 5 mg/ml, how many milliliters are required for a loading dose and three maintenance doses for a patient who weighs 220 pounds? (not 448.92, 149.85)

Chloroquine is used to prevent or suppress malaria in patients. Two weeks before the patient is potentially exposed to malaria, and for six weeks after the patient has returned, chloroquine is administered once a week (on the same day) at a dose of 500 mg. Suppose the patient spends 13 weeks in an area in which there is high exposure to malaria. How many 250-mg tablets of chloroquine will be administered altogether? (not 26)

Praziquantel is used to eradicate parasitic worms such as tapeworms. A one-day regimen consists of a dose of 60 mg for every kilogram of body mass. For a patient who weighs 154 pounds, how many 600-mg tablets of praziquantel are needed?
Answer #1

Metronidazole is used in the treatment of anaerobic protozoal infections. Although often administered orally, it can also be delivered by means of an IV infusion. The loading dose is 15 mg per kilogram of body mass and is infused over one hour. The maintenance dose is 7.5 mg per kilogram of body mass, is also infused over an hour, and is administered six hours after the loading dose. If the concentration of the metronidazole solution is 5 mg/ml, how many milliliters are required for a loading dose and three maintenance doses for a patient who weighs 220 pounds? (not 448.92, 149.85)

loading dose 15mg /kg over one hour

maintanance dose 7.5mg per kg and after 6 hours after loading dose

concentration 5mg/ml

weight 220 pounds = 99.79kg

loading dose= 15 * 99.79 = 1496.85mg

concentration 5mg/ ml

so, for 1496.85mg requires 1496.85/5 ml= 299.37ml

maintanance dose = 7.5mg/kg = 7.5* 99.79 = 748.42mg

so the dose will be 748.42/5 ml = 149.68ml (in loading dose it is not mentioned whether it is the whole dose for a day, it it is like that it has to be divided by 4 and given 37.42ml at a time )

Chloroquine is used to prevent or suppress malaria in patients. Two weeks before the patient is potentially exposed to malaria, and for six weeks after the patient has returned, chloroquine is administered once a week (on the same day) at a dose of 500 mg. Suppose the patient spends 13 weeks in an area in which there is high exposure to malaria. How many 250-mg tablets of chloroquine will be administered altogether? (not 26)

tablet has to be given two weeks before and during the expossure and 6 weeks after the return

so here 2+13+6 = 21 weeks the tablet should be given

dose is 500mg per week, so for 21 week, the number of 250mg table= 21*2 = 42 tablets

Praziquantel is used to eradicate parasitic worms such as tapeworms. A one-day regimen consists of a dose of 60 mg for every kilogram of body mass. For a patient who weighs 154 pounds, how many 600-mg tablets of praziquantel are needed?

dose= 60mg/kg

weight 154 pounds = 69.85kg

so dose for the patient = 60 * 69.85 = 4191mg

so, the 600mg tablet needed= 4191/600 = 6.985 aproximately 7 tablet required